# Reinforced Generation of Combinatorial Structures: Hardness of Approximation

### v5

@@ Appendix B Proofs of Average Case Theorems
    Soundness case: We must show that if $\max_{\mathbf{x}}I(\mathbf{x})\leq(1/k+\varepsilon)\cdot m$, then $\max_{\mathbf{x},\mathbf{y},\mathbf{z}}J(\mathbf{x},\mathbf{y},\mathbf{z})\leq(b+\varepsilon^{\prime})\cdot M$. Let $(\mathbf{x},\mathbf{y},\mathbf{z})$ be an assignment to the variables of $J$.
    The first step we perform is to assume without loss of generality that $\mathbf{y}\in Y$. To do this, note that by definition of $Y$, there is some permutation $\sigma:\Psi\to\Psi$ such that $\mathbf{y}^{\sigma}=(\sigma(y_{1}),\ldots,\sigma(y_{k}))\in Y$. We will apply $\sigma$ to all variables in the full assignment $(\mathbf{x},\mathbf{y},\mathbf{z})$, resulting in no change in the value of $J(\mathbf{x},\mathbf{y},\mathbf{z})$ while guaranteeing that $\mathbf{y}\in Y$.
    For $i\in\Psi$, let $\alpha_{i}$ denote the fraction of $P^{=}_{i}$ constraints of $I$ that are satisfied by $\mathbf{x}$. By assumption, we have $\sum_{i\in\Psi}\alpha_{i}\leq 1+k\varepsilon$. Let us define $\hat{\alpha_{i}}=\alpha_{i}/(\sum_{i^{\prime}\in\Psi}\alpha_{i^{\prime}})$, so $\sum_{i\in\Psi}\hat{\alpha_{i}}=1$. In other words, $\{\hat{\alpha_{i}}:i\in\Psi\}$ defines a probability distribution.
    On the other hand, since $y\in Y$ we can now bound $J(x,y,z)$ as follows.
    $J(\mathbf{x},\mathbf{y},\mathbf{z})$ $\leq m\cdot\left(\sum_{i\in\Psi}\alpha_{i}\cdot s_{1}(I_{i})+(1-\alpha_{i})\cdot s_{2}(I_{i})\right)$
    $\leq m\cdot\left(\sum_{i\in\Psi}\hat{\alpha_{i}}\cdot s_{1}(I_{i})+(1-\hat{\alpha_{i}})\cdot s_{2}(I_{i})\right)+O(\varepsilon)\cdot m$ ($\sum_{i\in\Psi}\alpha_{i}\leq 1+k\varepsilon$)
    $=m\cdot\left(\sum_{i\in\Psi}\hat{\alpha_{i}}\cdot(s_{1}(I_{i})-s_{2}(I_{i}))+s_{2}(I_{i})\right)+O(\varepsilon)\cdot m.$
    Note that subject to $\hat{\alpha_{i}}$ being a probability distribution, this is maximized when $\hat{\alpha_{i}}$ is a point mass on a maximizer of $s_{1}(I_{i})-s_{2}(I_{i})$. Therefore, this is bounded by
    $m\cdot\left(s_{1}(I_{r_{0}})+\sum_{i\in\Psi\setminus\{0\}}s_{2}(I_{r_{i}})\right)+O(\varepsilon)\cdot m=(b+O(\varepsilon))\cdot M=(b+\varepsilon^{\prime})\cdot M,$
@@ C.1 Description of our gadgets
For $k\in\{3,4\}$, we used AlphaEvolve to find gadgets $\{I_{i}:i\in\mathbb{Z}_{k}\}$, scoring a particular gadget by the inverse inapproximability ratio $a/b$, where $a$ and $b$ are defined in Theorem C.2. We now concretely define the specific gadgets found by AlphaEvolve and list the parameters they achieve (Definition C.2). We derive the final inapproximability results (Theorems 3.1 and 3.2) by applying Theorem C.2 on these gadgets.
We have formatted a gadget as a weighted “edge list” corresponding to a collection of $P^{\neq}$ clauses. The format is meant to be understood as follows: a tuple $(a,b,w)$ corresponds to $w$ parallel edges from variable $a$ to variable $b$, that is, the gadget contains $w$ copies of the term $P^{\neq}$ applied to variables $a$ and $b$.
@@
We write the list of edges for $I_{0}$ first, which has $12$ total variables, and hence $6$ auxiliary variables. Note that the global variables $\{4,5,6\}$ are not used in this gadget. See Figure 2(a) for a visual representation of the gadget.
@@
Below (and in Figure 2(b)), we write the list of edges for our $I_{1}$ gadget, which has $6+n_{\textup{aux}}=14$ total variables, and hence $8$ auxiliary variables. Note that repeated edges correspond to multiple $P^{\neq}$ clauses on the same pair of variables.
@@
$I_{2}$ is defined identically to $I_{1}$, except with the global variables reordered from $(4,5,6)$ to $(4,6,5)$. We can calculate the relevant parameters of these gadgets as below.
@@
Let $I_{0},I_{1}$, and $I_{2}$ be as above. $I_{0}$ is a 0-gadget from $3\textup{LIN}({3})$ to $\textup{MAX-}{3}\textup{-CUT}$ with parameters $c(I_{0})=s_{1}(I_{0})=t(I_{0})=18$, and $s_{2}(I_{0})=16$. For $i\in\{1,2\}$, $I_{i}$ is an $i$-gadget with parameters $c(I_{i})=s_{1}(I_{i})=48$, $s_{2}(I_{i})=46$, and $t(I_{i})=53$.
@@
Below we write the weighted edge list for our $I_{0}$ gadget for $\textup{MAX-}{4}\textup{-CUT}$.
@@
For $i\geq 1$, our $I_{i}$ gadget is obtained from our $I_{i-1}$ gadget by reordering the global variables from $(4,5,6,7)$ to $(7,4,5,6)$. The below Lemma can be verified computationally using Theorem C.1.
@@
Let $I_{0},I_{1}$, $I_{2}$, and $I_{3}$ be as above. For each $i$, $I_{i}$ is an $i$-gadget from $3\textup{LIN}({4})$ to $\textup{MAX-}{4}\textup{-CUT}$ with parameters $c(I_{i})=49535$, $s_{1}(I_{i})=49538$, $s_{2}(I_{i})=48681$, and $t(I_{i})=52941$.
@@ Appendix D Proof of the Hardness of Approximation for TSP
In this section we prove Theorem 4.1. We will begin by recalling the various predicates we consider in this section, all of which are over the boolean alphabet $\Psi=\{0,1\}$.
  • $P^{=}$ denotes a $2$-ary predicate defined by $P^{=}(x,y)=\mathbf{1}\{x=y\}$.
  • $P^{\neq}$ denotes a $2$-ary predicate defined by $P^{\neq}(x,y)=\mathbf{1}\{x\neq y\}$.
  • $P^{=}_{1}$ denotes a $3$-ary predicate defined by $P^{=}_{1}(x,y,z)=\mathbf{1}\{x+y+z\equiv 1\pmod{2}\}$.
@@
We will henceforth focus on proving inapproximability for $\textup{MWST}_{f}$ by reduction from $3\textup{LIN}({2})$. As in Section 4, our reduction is defined by a gadget that determines how we map $3\textup{LIN}({2})$ clauses to edges in a $\textup{MWST}_{f}$ instance, which we refer to as an equation gadget.
@@
Given an equation gadget $H$ and a valid spanning tour $Q$ in $H$, we define $k_{i}^{H}(Q)$ to be the number of special edges that appear exactly $i$ times in $Q$. Note for example that $\sum_{i}k_{i}^{H}(Q)=3$ for all valid $Q$.
$H$ will be used to map the “equals 1” clause $P^{=}_{1}(z_{1},z_{2},z_{3})=\mathbf{1}\{z_{1}+z_{2}+z_{3}\equiv 1\pmod{2}\}$ of $3\textup{LIN}({2})$ (recall that the variables $z_{\ell}$ are over $\{0,1\}$). As intuition, we will aim to encode an assignment $\mathbf{z}\in\{0,1\}^{3}$ to the variables of a $P^{=}_{1}$ clause as a spanning tour $T$ such that the special edge $(i,4)$ is used exactly $2\cdot z_{i}$ times. We will require that satisfying assignments to $P^{=}_{1}$ admit such spanning tours (which must have $k_{1}(T)=0$ and $k_{2}(T)$ odd) with low weight, and that spanning tours not of this form must be more expensive. We formally encapsulate this intuition in the following definition.
@@
  • Completeness: Let $\mathbf{z}\in\{0,1\}^{3}$, and let $C=(P_{1}^{=})^{-1}(1)\subset\{0,1\}^{3}$ be the set of satisfying assignments of $P_{1}^{=}$. We use $\mathbb{Q}_{\mathbf{z}}$ to refer to the set of valid spanning tours $Q$ in $H$ such that the special edge $(\ell,4)$ is unused for all $\ell$ such that $\mathbf{z}_{\ell}=0$, and is used twice otherwise. $c(H)$ is defined as
@@
Similarly to [11], we will perform a reduction from a particular “hybrid” bounded-occurrence CSP over $\{0,1\}$ (described in Theorem D.1). Informally, the hybrid CSP is a particularly structured instance that contains a small number of $P^{=}_{1}$ clauses along with some simpler two-variable equality and inequality clauses, where $P^{=}_{1}(z_{1},z_{2},z_{3})=\mathbf{1}\{z_{1}+z_{2}+z_{3}\equiv 1\pmod{2}\}$ defined earlier. Consider the 2-ary predicates $P^{=}(x_{1},x_{2})=\mathbf{1}\{x_{1}=x_{2}\}$, $P^{\neq}(x_{1},x_{2})=\mathbf{1}\{x_{1}\neq x_{2}\}$. In Definition D.5 we provide the CSP instance we reduce from, and in Theorem D.1 we state its soundness and completeness guarantees.
@@
Let $k$ be a fixed integer. An instance of $\text{Hybrid-}3\textup{LIN}({2})$ takes the form $I=I^{=}+2I^{\neq}+2I_{1}^{=}$ of $\textup{CSP}(P_{1}^{=},P^{=},P^{\neq})$ where:
  • Variables: The variables in $I$ are labeled as $x_{i,j,b}$ for $i\in[n]$, $j\in[11k]$, and $b\in\{0,1\}$. That is, there are $22kn$ variables. Each variable will participate in exactly two distinct $P^{=}$ clauses in $I^{=}$. Let $S=\{11\ell:\ell\in[k]\}$ be the set of multiples of $11$ in $[11k]$. Following the notation in [33], we will classify a variable $x_{i,j,b}$ as a “contact” if $j\in S$, and a “checker” otherwise. A contact will additionally appear in exactly one $P^{\neq}$ clause in $I^{\neq}$, and a checker will appear in exactly one $P_{1}^{=}$ clause in $I_{1}^{=}$.
  • Equality clauses: $I^{=}$ consists of the following: each variable $x_{i,j,b}$ participates in two clauses $P^{=}(x_{i,j,b},x_{i,j+1,b})$ and $P^{=}(x_{i,j,b},x_{i,j-1,b})$, where addition/subtraction to $j$ is done modulo $11k$.
  • Inequality clauses: $I^{\neq}$ is defined by the following: for each $i\in[n]$, $j\in[11k]\setminus S$, we add the clause $P^{\neq}(x_{i,j,0},x_{i,M(j),1})$, where $M$ is a fixed bijection on $11k]\setminus S$.
  • 3LIN(2) clauses: $I_{1}^{=}$ consists of a collection of $P_{1}^{=}$ clauses involving only contact variables, and every contact variable participates in exactly one such clause.
We remark that $3\textup{LIN}({2})$ technically refers to a CSP rather than a clause in a CSP. However we will (as we did in Definition D.5 and in the discussion before) refer to $P^{=}_{1}$ clauses as $3\textup{LIN}({2})$ clauses” for better readability. Note that in the definition above, $I$ contains one copy of each clause from $I^{=}$, and two copies of each clause from $I^{\neq}$ and $I_{1}^{=}$.
@@
    Completeness: There is an assignment $\mathbf{x}=(x_{i,j,b})_{i,j,b}$ satisfying all clauses in $I^{=}$ and $I^{\neq}$, and all but $\varepsilon nk$ clauses in $I_{1}^{=}$, or
@@
    Soundness: For all assignments $\mathbf{x}=(x_{i,j,b})_{i,j,b}$, at least $(2/3-\varepsilon)nk$ clauses of $I=I^{=}+2I^{\neq}+2I_{1}^{=}$ are unsatisfied.
@@
We will provide a reduction from the NP-hard problem described in Theorem D.1 to $\textup{MWST}_{f}$. Suppose we are given a $\text{Hybrid-}3\textup{LIN}({2})$ instance $I=I^{=}+2I^{\neq}+2I_{1}^{=}$ on variables $x_{i,j,b}$ for $i\in[n],j\in[11k],b\in\{0,1\}$. Our instance $G$ of $\textup{MWST}_{f}$ will contain a vertex for each variable, which we will also denote by $x_{i,j,b}$. Additionally we will add a single “central” vertex $s$. We will convert every clause of $I$ into a corresponding structure in $G$, which we define below. We also provide a visual representation of this correspondence in Figure 7(a) and Figure 7(b).
  • Equality clauses: Each clause in $I^{=}$ will be replaced by a single unforced edge in $G$ of weight one. (In Figure 7(b) these edges are denoted in black.)
  • Inequality clauses: Each clause in $I^{\neq}$ will be replaced by two parallel forced edges in $G$ of weight one. (In Figure 7(b) these edges are denoted in red.)
  • 3LIN(2) clauses: A clause $P^{=}_{1}(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})$ in $I_{1}^{=}$ will be replaced by a copy of the gadget $H$ (shown in Figure 7(a)). More precisely, for each such clause we will add $n_{\textup{aux}}$ new vertices, say $a_{1},\ldots,a_{n_{\textup{aux}}}$, and add a single induced copy of the forced and unforced edges from $H$ on the contact vertices $x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}}$, auxiliary vertices $a_{1},\ldots,a_{n_{\textup{aux}}}$, and central vertex $s$. Note that the central vertex $s$ is shared among all $P^{=}_{1}$ clauses, but the auxiliary vertices are individually created for each $P^{=}_{1}$ clause.
We will define $^{=}, EE^{\neq}$, and $E_{1}^{=}$ as the edges added in the three steps above respectively. This completes the definition of $G$. It remains to argue that the $\textup{MWST}_{f}$ value of $G$ has the desired gap in the completeness and soundness cases.
@@
(b) Graph encoding $I^{=}$ and $I^{\neq}$ clauses. The green vertices correspond to the contact vertices, and the white ones correspond to the checker vertices. The black edges encode the $I^{=}$ clauses, and the red edges encode the $I^{\neq}$ clauses. All edges have weight one.
@@
Suppose there is an assignment $\mathbf{x}=(x_{i,j,b})_{i,j,b}$ to the $\text{Hybrid-}3\textup{LIN}({2})$ instance $I$ that satisfies all clauses in $I^{=}$ and $I^{\neq}$ clauses, and violates at most $\Delta$ clauses in $I_{1}^{=}$. Then, there is a valid spanning tour in $G$ of total weight at most $nk\cdot(30+2\cdot c(H)/3)+2(\Delta+2n)\cdot w(H)$.

### v6

@@ Appendix B Proofs of Average Case Theorems
    Soundness case: We must show that if $\max_{\mathbf{x}}\texttt{I}(\mathbf{x})\leq(1/k+\varepsilon)\cdot m$, then $\max_{\mathbf{x},\mathbf{y},\mathbf{z}}\texttt{J}(\mathbf{x},\mathbf{y},\mathbf{z})\leq(b+\varepsilon^{\prime})\cdot M$. Let $(\mathbf{x},\mathbf{y},\mathbf{z})$ be an assignment to the variables of J.
    The first step we perform is to assume without loss of generality that $\mathbf{y}\in Y$. To do this, note that by definition of $Y$, there is some permutation $\sigma:\mathbb{Z}_{k}\to\mathbb{Z}_{k}$ such that $\mathbf{y}^{\sigma}=(\sigma(y_{1}),\ldots,\sigma(y_{k}))\in Y$. We will apply $\sigma$ to all variables in the full assignment $(\mathbf{x},\mathbf{y},\mathbf{z})$, resulting in no change in the value of $\texttt{J}(\mathbf{x},\mathbf{y},\mathbf{z})$ while guaranteeing that $\mathbf{y}\in Y$.
    For $i\in\mathbb{Z}_{k}$, let $\alpha_{i}$ denote the fraction of $\texttt{P}^{\equiv i}_{3,k}$ constraints of $I$ that are satisfied by $\mathbf{x}$. By assumption, we have $\sum_{i\in\mathbb{Z}_{k}}\alpha_{i}\leq 1+k\varepsilon$. Let us define $\hat{\alpha_{i}}=\alpha_{i}/(\sum_{i^{\prime}\in\mathbb{Z}_{k}}\alpha_{i^{\prime}})$, so $\sum_{i\in\mathbb{Z}_{k}}\hat{\alpha_{i}}=1$. In other words, $\{\hat{\alpha_{i}}:i\in\mathbb{Z}_{k}\}$ defines a probability distribution.
    On the other hand, since $y\in Y$ we can now bound $\texttt{J}(\mathbf{x},\mathbf{y},\mathbf{z})$ as follows.
    $\texttt{J}(\mathbf{x},\mathbf{y},\mathbf{z})$ $\leq m\cdot\left(\sum_{i\in\mathbb{Z}_{k}}\alpha_{i}\cdot s_{1}(\texttt{I}^{\equiv i}_{3,k})+(1-\alpha_{i})\cdot s_{2}(\texttt{I}^{\equiv i}_{3,k})\right)$
    $\leq m\cdot\left(\sum_{i\in\mathbb{Z}_{k}}\hat{\alpha_{i}}\cdot s_{1}(\texttt{I}^{\equiv i}_{3,k})+(1-\hat{\alpha_{i}})\cdot s_{2}(\texttt{I}^{\equiv i}_{3,k})\right)+O(\varepsilon)\cdot m$ ($\sum_{i\in\mathbb{Z}_{k}}\alpha_{i}\leq 1+k\varepsilon$)
    $=m\cdot\left(\sum_{i\in\mathbb{Z}_{k}}\hat{\alpha_{i}}\cdot(s_{1}(\texttt{I}^{\equiv i}_{3,k})-s_{2}(\texttt{I}^{\equiv i}_{3,k}))+s_{2}(\texttt{I}^{\equiv i}_{3,k})\right)+O(\varepsilon)\cdot m.$
    Note that subject to $\hat{\alpha_{i}}$ being a probability distribution, this is maximized when $\hat{\alpha_{i}}$ is a point mass on a maximizer of $s_{1}(\texttt{I}^{\equiv i}_{3,k})-s_{2}(\texttt{I}^{\equiv i}_{3,k})$. Therefore, this is bounded by
    $m\cdot\left(s_{1}(\texttt{I}^{\equiv r_{0}}_{+,)}\sum_{i\in\mathbb{Z}_{k}\setminus\{0\}}s_{2}(\texttt{I}^{\equiv r_{i}}_{3,k})\right)+O(\varepsilon)\cdot m=(b+O(\varepsilon))\cdot M=(b+\varepsilon^{\prime})\cdot M,$
@@ C.1 Description of our gadgets
For $k\in\{3,4\}$, we used AlphaEvolve to find gadgets $\{\texttt{I}^{\equiv i}_{3,k}:i\in\mathbb{Z}_{k}\}$, scoring a particular gadget by the inverse inapproximability ratio $a/b$, where $a$ and $b$ are defined in Theorem C.2. We now concretely define the specific gadgets found by AlphaEvolve and list the parameters they achieve (Definition C.2). We derive the final inapproximability results (Theorems 3.1 and 3.2) by applying Theorem C.2 on these gadgets.
We have formatted a gadget as a weighted “edge list” corresponding to a collection of $\texttt{P}^{\neq}_{2}$ clauses. The format is meant to be understood as follows: a tuple $(a,b,w)$ corresponds to $w$ parallel edges from variable $a$ to variable $b$, that is, the gadget contains $w$ copies of the term $\texttt{P}^{\neq}_{2}$ applied to variables $a$ and $b$.
@@
We write the list of edges for $\texttt{I}^{\equiv 0}_{3,3}$ first, which has $12$ total variables, and hence $6$ auxiliary variables. Note that the global variables $\{4,5,6\}$ are not used in this gadget. See Figure 2(a) for a visual representation of the gadget.
@@
Below (and in Figure 2(b)), we write the list of edges for our $\texttt{I}^{\equiv 1}_{3,3}$ gadget, which has $6+n_{\textup{aux}}=14$ total variables, and hence $8$ auxiliary variables. Note that repeated edges correspond to multiple $\texttt{P}^{\neq}_{2}$ clauses on the same pair of variables.
@@
$\texttt{I}^{\equiv 2}_{3,3}$ is defined identically to $\texttt{I}^{\equiv 1}_{3,3}$, except with the global variables reordered from $(4,5,6)$ to $(4,6,5)$. We can calculate the relevant parameters of these gadgets as below.
@@
Let $\texttt{I}^{\equiv 0}_{3,3},\texttt{I}^{\equiv 1}_{3,3}$, and $\texttt{I}^{\equiv 2}_{3,3}$ be as above. $\texttt{I}^{\equiv 0}_{3,3}$ is a 0-gadget that maps the predicate $\texttt{P}^{\equiv 0}_{3,3}$ to an instance of $\textup{MAX-}{3}\textup{-CUT}$, with parameters $c(\texttt{I}^{\equiv 0}_{3,3})=s_{1}(\texttt{I}^{\equiv 0}_{3,3})=t(\texttt{I}^{\equiv 0}_{3,3})=18$, and $s_{2}(\texttt{I}^{\equiv 0}_{3,3})=16$. For $i\in\{1,2\}$, $\texttt{I}^{\equiv i}_{3,3}$ is an $i$-gadget with parameters $c(\texttt{I}^{\equiv i}_{3,3})=s_{1}(\texttt{I}^{\equiv i}_{3,3})=48$, $s_{2}(\texttt{I}^{\equiv i}_{3,3})=46$, and $\texttt{count}(\texttt{I}^{\equiv i}_{3,3})=53$.
@@
Below we write the weighted edge list for our $\texttt{I}^{\equiv 0}_{3,4}$ gadget for $\textup{MAX-}{4}\textup{-CUT}$.
@@
For $i\geq 1$, our $\texttt{I}^{\equiv i}_{3,4}$ gadget is obtained from our $\texttt{I}^{\equiv(i-1)}_{3,4}$ gadget by reordering the global variables from $(4,5,6,7)$ to $(7,4,5,6)$. The below Lemma can be verified computationally using Theorem C.1.
@@
Let $\texttt{I}^{\equiv 0}_{3,4},\texttt{I}^{\equiv 1}_{3,4}$, $\texttt{I}^{\equiv 2}_{3,4}$, and $\texttt{I}^{\equiv 3}_{3,4}$ be as above. For each $i$, $\texttt{I}^{\equiv i}_{3,4}$ is a $i$-gadget that maps the predicate $\texttt{P}^{\equiv i}_{3,4}$ to an instance of $\textup{MAX-}{4}\textup{-CUT}$, with parameters $c(\texttt{I}^{\equiv i}_{3,4})=49535$, $s_{1}(\texttt{I}^{\equiv i}_{3,4})=49538$, $s_{2}(\texttt{I}^{\equiv i}_{3,4})=48681$, and $\texttt{count}(\texttt{I}^{\equiv i}_{3,4})=52941$.
@@ Appendix D Proof of the Hardness of Approximation for TSP
In this section we prove Theorem 4.1. We will begin by recalling the various predicates we consider in this section, all of which are over the boolean alphabet $\mathbb{Z}_{2}=\{0,1\}$. Following the convention in the rest of the paper, we define the predicates relevant to this section.
  • $\texttt{P}^{=}_{2}$ denotes a binary predicate defined by $\texttt{P}^{=}_{2}(x,y)=\mathbf{1}\{x=y\}$.
  • $\texttt{P}^{\neq}_{2}$ denotes a binary predicate defined by $\texttt{P}^{\neq}_{2}(x,y)=\mathbf{1}\{x\neq y\}$.
  • $\texttt{P}^{\equiv 1}_{3,2}$ denotes a $3$-ary predicate defined by $\texttt{P}^{\equiv 1}_{3,2}(x,y,z)=\mathbf{1}\{x+y+z\equiv 1\pmod{2}\}$.
@@
We will henceforth focus on proving inapproximability for $\textup{MWST}_{f}$ by reduction from $3\textup{LIN}({2})$. As in Section 4, our reduction is defined by a gadget that determines how we map $\texttt{P}^{\equiv 1}_{3,2}$ clauses to edges in a $\textup{MWST}_{f}$ instance, which we refer to as an equation gadget.
@@
Given an equation gadget $H$ and a valid spanning tour $Q$ in $H$, we define
  • $k_{i}^{H}(Q)$ to be the number of special edges that appear exactly I times in $Q$. Note for example that $\sum_{i}k_{i}^{H}(Q)=3$ for all valid $Q$.
$H$ will be used to map the “equals 1” clause $\texttt{P}^{\equiv 1}_{3,2}(z_{1},z_{2},z_{3})=\mathbf{1}\{z_{1}+z_{2}+z_{3}\equiv 1\pmod{2}\}$ of $3\textup{LIN}({2})$ (recall that the variables $z_{\ell}$ are over $\{0,1\}$). As intuition, we will aim to encode an assignment $\mathbf{z}\in\{0,1\}^{3}$ to the variables of a $\texttt{P}^{\equiv 1}_{3,2}$ clause as a spanning tour $T$ such that the special edge $(i,4)$ is used exactly $2\cdot z_{i}$ times. We will require that satisfying assignments to $\texttt{P}^{\equiv 1}_{3,2}$ admit such spanning tours (which must have $k_{1}(T)=0$ and $k_{2}(T)$ odd) with low weight, and that spanning tours not of this form must be more expensive. We formally encapsulate this intuition in the following definition.
@@
  • Completeness: Let $\mathbf{z}\in\{0,1\}^{3}$, and let $C=(\texttt{P}^{\equiv 1}_{3,2})^{-1}(1)\subset\{0,1\}^{3}$ be the set of satisfying assignments of $\texttt{P}^{\equiv 1}_{3,2}$. We use $\mathbb{Q}_{\mathbf{z}}$ to refer to the set of valid spanning tours $Q$ in $H$ such that the special edge $(\ell,4)$ is unused for all $\ell$ such that $\mathbf{z}_{\ell}=0$, and is used twice otherwise. $c(H)$ is defined as
@@
Similarly to [15], we will perform a reduction from a particular “hybrid” bounded-occurrence CSP over $\{0,1\}$ (described in Theorem D.1). Informally, the hybrid CSP is a particularly structured instance that contains a small number of $\texttt{P}^{\equiv 1}_{3,2}$ clauses along with some simpler two-variable equality and inequality clauses, where $\texttt{P}^{\equiv 1}_{3,2}(z_{1},z_{2},z_{3})=\mathbf{1}\{z_{1}+z_{2}+z_{3}\equiv 1\pmod{2}\}$. Consider the 2-ary predicates $\texttt{P}^{=}_{2}(x_{1},x_{2})=\mathbf{1}\{x_{1}=x_{2}\}$, $\texttt{P}^{\neq}_{2}(x_{1},x_{2})=\mathbf{1}\{x_{1}\neq x_{2}\}$. In Definition D.5 we provide the CSP instance we reduce from, and in Theorem D.1 we state its soundness and completeness guarantees.
@@
Let $k$ be a fixed integer. An instance of $\text{Hybrid-}3\textup{LIN}({2})$ takes the form $\texttt{I}=\texttt{I}^{=}_{2}+2\cdot\texttt{I}^{\neq}_{2}+2\cdot\texttt{I}^{\equiv 1}_{3,2}$ of $\textup{CSP}(\texttt{P}^{\equiv 1}_{3,2},\texttt{P}^{=}_{2},\texttt{P}^{\neq}_{2})$ where:
  • Variables: The variables in I are labeled as $x_{i,j,b}$ for $i\in[n]$, $j\in[11k]$, and $b\in\{0,1\}$. That is, there are $22kn$ variables. Each variable will participate in exactly two distinct $\texttt{P}^{=}_{2}$ clauses in $\texttt{I}^{=}_{2}$. Let $S=\{11\ell:\ell\in[k]\}$ be the set of multiples of $11$ in $[11k]$. Following the notation in [39], we will classify a variable $x_{i,j,b}$ as a “contact” if $j\in S$, and a “checker” otherwise. A contact will additionally appear in exactly one $\texttt{P}^{\neq}_{2}$ clause in $\texttt{I}^{\neq}_{2}$, and a checker will appear in exactly one $\texttt{P}^{\equiv 1}_{3,2}$ clause in $\texttt{I}^{\equiv 1}_{3,2}$.
  • Equality clauses: $\texttt{I}^{=}_{2}$ consists of the following: each variable $x_{i,j,b}$ participates in two clauses $\texttt{P}^{=}_{2}(x_{i,j,b},x_{i,j+1,b})$ and $\texttt{P}^{=}_{2}(x_{i,j,b},x_{i,j-1,b})$, where addition/subtraction to $j$ is done modulo $11k$.
  • Inequality clauses: $\texttt{I}^{\neq}_{2}$ is defined by the following: for each $i\in[n]$, $j\in[11k]\setminus S$, we add the clause $\texttt{P}^{\neq}_{2}(x_{i,j,0},x_{i,\pi(j),1})$, where $\pi$ is a fixed bijection on $11k]\setminus S$.
  • 3LIN(2) clauses: $\texttt{I}^{\equiv 1}_{3,2}$ consists of a collection of $\texttt{P}^{\equiv 1}_{3,2}$ clauses involving only contact variables, and every contact variable participates in exactly one such clause.
Note that in the definition above, I contains one copy of each clause from $\texttt{I}^{=}_{2}$, and two copies of each clause from $\texttt{I}^{\neq}_{2}$ and $\texttt{I}^{\equiv 1}_{3,2}$.
@@
    Completeness: There is an assignment $\mathbf{x}=(x_{i,j,b})_{i,j,b}$ satisfying all clauses in $\texttt{I}^{=}_{2}$ and $\texttt{I}^{\neq}_{2}$, and all but $\varepsilon nk$ clauses in $\texttt{I}^{\equiv 1}_{3,2}$, or
@@
    Soundness: For all assignments $\mathbf{x}=(x_{i,j,b})_{i,j,b}$, at least $(2/3-\varepsilon)nk$ clauses of $\texttt{I}=\texttt{I}^{=}_{2}+2\texttt{I}^{\neq}_{2}+2\texttt{I}^{\equiv 1}_{3,2}$ are unsatisfied.
@@
We will provide a reduction from the NP-hard problem described in Theorem D.1 to $\textup{MWST}_{f}$. Suppose we are given a $\text{Hybrid-}3\textup{LIN}({2})$ instance $\texttt{I}=\texttt{I}^{=}_{2}+2\cdot\texttt{I}^{\neq}_{2}+2\cdot\texttt{I}^{\equiv 1}_{3,2}$ on variables $x_{i,j,b}$ for $i\in[n],j\in[11k],b\in\{0,1\}$. Our instance $G$ of $\textup{MWST}_{f}$ will contain a vertex for each variable, which we will also denote by $x_{i,j,b}$. Additionally we will add a single “central” vertex $s$. We will convert every clause of I into a corresponding structure in $G$, which we define below. We also provide a visual representation of this correspondence in Figure 7(a) and Figure 7(b).
  • Equality clauses: Each clause in $\texttt{I}^{=}_{2}$ will be replaced by a single unforced edge in $G$ of weight one. (In Figure 7(b) these edges are denoted in black.)
  • Inequality clauses: Each clause in $\texttt{I}^{\neq}_{2}$ will be replaced by two parallel forced edges in $G$ of weight one. (In Figure 7(b) these edges are denoted in red.)
  • 3LIN(2) clauses: A clause $\texttt{P}^{\equiv 1}_{3,2}(x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}})$ in $\texttt{I}^{\equiv 1}_{3,2}$ will be replaced by a copy of the gadget $H$ (shown in Figure 7(a)). More precisely, for each such clause we will add $n_{\textup{aux}}$ new vertices, say $a_{1},\ldots,a_{n_{\textup{aux}}}$, and add a single induced copy of the forced and unforced edges from $H$ on the contact vertices $x_{i_{1},j_{1},b_{1}},x_{i_{2},j_{2},b_{2}},x_{i_{3},j_{3},b_{3}}$, auxiliary vertices $a_{1},\ldots,a_{n_{\textup{aux}}}$, and central vertex $s$. Note that the central vertex $s$ is shared among all $\texttt{P}^{\equiv 1}_{3,2}$ clauses, but the auxiliary vertices are individually created for each $\texttt{P}^{\equiv 1}_{3,2}$ clause.
We will define $\texttt{E}^{=}_{2}$, $\texttt{E}^{\neq}_{2}$, and $\texttt{E}^{\equiv 1}_{3,2}$ as the edges added in the three steps above respectively. This completes the definition of $G$. It remains to argue that the $\textup{MWST}_{f}$ value of $G$ has the desired gap in the completeness and soundness cases.
@@
(b) Graph encoding $\texttt{I}^{=}_{2}$ and $\texttt{I}^{\neq}_{2}$ clauses. The green vertices correspond to the contact vertices, and the white ones correspond to the checker vertices. The black edges encode the $\texttt{I}^{=}_{2}$ clauses, and the red edges encode the $\texttt{I}^{\neq}_{2}$ clauses. All edges have weight one.
@@
Suppose there is an assignment $\mathbf{x}=(x_{i,j,b})_{i,j,b}$ to the $\text{Hybrid-}3\textup{LIN}({2})$ instance I that satisfies all clauses in $\texttt{I}^{=}_{2}$ and $\texttt{I}^{\neq}_{2}$ clauses, and violates at most $\Delta$ clauses in $\texttt{I}^{\equiv 1}_{3,2}$. Then, there is a valid spanning tour in $G$ of total weight at most $nk\cdot(30+2\cdot c(H)/3)+2(\Delta+2n)\cdot w(H)$.